BitBox has raised S10 million in a Series A round with S40 million post-money value and a 1.5x liquidation preference, and S25 million in a Series B round with a S75 million post-money value and a 3x liquidation preference plus seniority over Series A. What will Series A, Series B, and common shareholders receive if BitBox is sold for a. $85 million? b. S100 million? c. S200 million? d. S300 million?
Answer #1

Series A liquidation preference = 10 * 1.5 = 15 million,

Ownership = 10/40 * 50/75 = 16.67%,

Series B liquidation preference = 25 * 3 = 75 million,

Ownership = 25/75 = 33.33%.

Common owns 50%.

Part A: $75 to B, $10 to A.

Part B: $75 to B, $15 to A, $10 to common.

Part C: 75 to B, A converts and gets $125 * (16.67/ (16.67 + 50)) = $31.25 million, common gets $93.75.

Part D: A and B convert, so A gets 16.67% * (300) = 50,

B gets 33.3% * (300) = 100, and common gets $150.